OpenStudy (anonymous):

what is the slope of -3x+4y=5

OpenStudy (midhun.madhu1987):

Convert into the slope intercept form: y = mx + c where m = slope and c = y-intercept...

OpenStudy (precal):

do you know how to do this?

OpenStudy (anonymous):

take \[-3x \] to the RHS, and divide by constant with \[y\] then \[y = \frac{3x}{4}+ \frac{5}{4} , \] then slope = \[\frac{3}{4}\]